OpenStudy (anonymous):

What is the value of 10C7? A.12 B.720 C.5040 D.604,800

OpenStudy (anonymous):

10C7=10C3=(10*9*8)/6=12

OpenStudy (anonymous):

120

OpenStudy (anonymous):

ok thanks i messed up 120 is the first one thanks :)

OpenStudy (anonymous):

because nCr=nC(n-r)...i used this property ,do you get it???

OpenStudy (anonymous):

ya i think so thanks